OpenStudy (anonymous):

Find the sum of the measures of the interior angles of a convex 20-gon.

OpenStudy (anonymous):

Sum = (n-2)*180

OpenStudy (anonymous):

\[Sum = (n-2) \times 180 = 18 \times 180 = 3240\]

OpenStudy (anonymous):

n is 20 here..